Question on Arithmetic Progression (AP)
Find the S15 if T4 = 15 and T9 = 30

Answer by ikleyn(52788)

.
Question on Arithmetic Progression (AP)
Find the S15 if T4 = 15 and T9 = 30
~~~~~~~~~~~~~~~~~~~~~~~~~~

You are given 

T4 = 15 = T1 + 3*d,   (1)
T9 = 30 = T1 + 8*d.   (2)

Distract (1) from (2) and get 5d = 15,  d = 3.

Then from (1) find T1 = 15 - 3*d = 15 - 9 = 6.

So, your AP has the first term T1 = 6 and the common difference 3.

Now apply the formula for the sum of AP.

For arithmetic progressions, see the lessons
    - Arithmetic progressions
    - The proofs of the formulas for arithmetic progressions
    - Problems on arithmetic progressions
    - Word problems on arithmetic progressions
    - Chocolate bars and arithmetic progressions
    - Mathematical induction and arithmetic progressions
    - One characteristic property of arithmetic progressions
    - Mathematical induction for sequences other than arithmetic or geometric
    - Proving inequalities by the method of Mathematical Induction
    - Solved problems on arithmetic progressions

OVERVIEW of these lessons on arithmetic progressions with short annotations is in the lesson  OVERVIEW of lessons on arithmetic progressions.